# 3-manifolds with universal links are spherical

### v1

@@ 6. Further comments
Berestein and Edmons proved in [BeresteinEdmonds:BranchedCovers], among other things, that an orientable $3$-manifold $M$ is a $3$-fold branched covering of $\mathbb{S}^{2}\times\mathbb{S}^{1}$ if and only if $H^{1}(M)\neq 0$. Some time later, Núñez in [Nuñez:UniversalLinksInS2XS1] gave an example of a link $L$ in $\mathbb{S}^{2}\times\mathbb{S}^{1}$ such that every closed, orientable $3$-manifold NN with $H^{1}(N)\neq 0$ is a branched covering over $\mathbb{S}^{2}\times\mathbb{S}^{1}$ with branching set exactly the link $L,$ and we denote these links as universal link for $\mathbb{S}^{2}\times\mathbb{S}^{1}.$ In the case of aspherical $3$-manifolds we ask:
Question 6.1.
Does there exist a closed, aspherical, orientable $3$-manifold $M$ and a link $L$ inside, such that every closed, aspherical, orientable $3$-manifold is a branched covering over $M$ with branching set exactly the link $L$?
And the next problem:
Problem 6.2.
@@
As we saw in Theorem 5.3, the notions of universal link and universal branching link are not equivalent for non-spherical $3$-manifolds, but one could ask whether these notions coincide in the case of spherical $3$-manifolds.
Question 6.3.

### v3

@@ 5. Further comments
As we point out in Section 4, Núñez gave in [Nuñez:UniversalLinksInS2XS1] an example of a universal link for $\mathbb{S}^{2}\times\mathbb{S}^{1}$ (see Definition 4.1). It will be interesting to have a more complete result:
Problem 5.1.
@@
Notice that for some other topological constrains, e.g. being non-orientable, the analogue of Proposition 4.2 and Corollary 4.4 is false, because Berestein and Edmonds proved in [BeresteinEdmonds:BranchedCovers] that every closed non-orientable $3$-manifold is a branched covering of $P^{2}\times\mathbb{S}^{1},$ but as pointed in [Nuñez:UniversalLinksInS2XS1] there is no example of a universal link for $P^{2}\times\mathbb{S}^{1}.$ Then we have the following question:
Problem 5.2.
Decide if there exist a link $L$ inside $P^{2}\times\mathbb{S}^{1},$ such that every closed, non-orientable $3$-manifold NN is a branched covering over $P^{2}\times\mathbb{S}^{1}$ with branching set exactly the link $L.$
Following the opposite direction of Theorem 4.3, we have that $\mathbb{S}^{2}\times\mathbb{S}^{1}$ is a $2$-folds cover $P^{3}\#P^{3},$ then every closed, orientable $3$-manifold NN with $H^{1}(N)\neq 0$ is a branched cover over $P^{3}\#P^{3}.$ Then we could ask:
Problem 5.3.
Identify which closed $3$-manifold branch cover $P^{3}\#P^{3}.$
In the topic of complement universal link, we saw in Corollary 3.3, the notions of universal link and universal branching link are not equivalent for non-spherical $3$-manifolds, but one could ask whether these notions coincide in the case of spherical $3$-manifolds.
Question 5.4.